OpenStudy (anonymous):

Determine whether the graphs of y=2x+4 and x+2y=6 are parallel lines

OpenStudy (anonymous):

Look at the slope (number next to x). Since it is the same number in both it is parallel.

OpenStudy (anonymous):

m=2 in first equation and 2y=-x+6 so y=-1/2x+3 so m=-1/2 so slopes are not same not parallel (they are actually perpendicular)

OpenStudy (anonymous):

Yeah, the question is trying to trick you. First you have to solve for y, the look at it.

OpenStudy (anonymous):

oops it's not the same.. so they are not parallel. I took y for x :S

OpenStudy (anonymous):

Its okay and thank you everyone